# MMBind: Unleashing the Potential of Distributed and Heterogeneous Data for Multimodal Learning in IoT

### v1

@@ Abstract.
Multimodal sensing systems are increasingly prevalent in various real-world applications. Most existing multimodal learning approaches heavily rely on training with a large amount of complete multimodal data. However, such a setting is impractical in real-world IoT sensing applications where data is typically collected by distributed nodes with heterogeneous data modalities, and is also rarely labeled. In this paper, we propose MMBind, a new framework for multimodal learning on distributed and heterogeneous IoT data. The key idea of MMBind is to construct a pseudo-paired multimodal dataset for model training by binding data from disparate sources and incomplete modalities through a sufficiently descriptive shared modality. We demonstrate that data of different modalities observing similar events, even captured at different times and locations, can be effectively used for multimodal training. Moreover, we propose an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations, coupled with a weighted contrastive learning approach to handle domain shifts among disparate data. Evaluations on ten real-world multimodal datasets highlight that MMBind outperforms state-of-the-art baselines under varying data incompleteness and domain shift, and holds promise for advancing multimodal foundation model training in IoT applications [^1] The source code for MMBind will be released once the paper is accepted..
copyright: nonedoi: isbn:
@@ 1. Introduction
Multimodal sensing systems are increasingly deployed in real-world applications such as health monitoring (Ouyang et al., 2024; Zhang et al., 2020), Augmented Reality (AR) (Chen et al., 2018; Apicharttrisorn et al., 2022), localization (Jeong et al., 2024), and autonomous driving (Shi et al., 2022; He et al., 2023). The complexity and dynamic nature of these tasks often surpass the capabilities of a single sensor modality, requiring the integration of multiple complementary data modalities to enhance performance.
Most existing multimodal learning approaches require extensive training with complete multimodal data, where each sample includes data from all sensor modalities collected simultaneously along with corresponding labels (Ouyang et al., 2022; Liu et al., 2024, 2021). However, this paradigm is impractical in real-world IoT sensing applications where data is typically collected by distributed nodes. First, such data is usually heterogeneous and incomplete—the available multimodal data on different nodes can vary significantly, often with missing modalities or labels. For example, sensor nodes deployed in homes may contain diverse sets of modalities due to deployment constraints or privacy concerns. Sensors on a specific node may also fail dynamically, resulting in changing sensor modalities at runtime. Labeling multimodal data in real-world settings is also challenging, as data from many sensors, such as IMU, are not intuitive for human annotation (Hu et al., 2020; Shuai et al., 2021). Second, sensor data collected by distributed nodes is typically gathered at different times and locations, resulting in disparate data that may describe similar events. For instance, in activity recognition, a wearable worn by an outdoor subject may only collect IMU data, while an indoor node may exclusively capture skeleton data of the same activity. Given the prevalence of disparate and heterogeneous data in IoT sensing, it is vital to effectively incorporate this data for multimodal learning.
Figure 1. Conventional multimodal learning requires complete multimodal data. MMBind features a data binding approach that effectively integrates disparate data with heterogeneous modalities for multimodal learning, which also outperforms the model binding approach, where various modalities are indirectly aligned through the encoder of a shared modality.
@@
In this paper, we propose MMBind, a new framework for multimodal learning with distributed and heterogeneous IoT data. As shown in Figure 1, the key idea of MMBind is to leverage shared modalities to bind disparate and incomplete data collected by distributed nodes, thereby creating pseudo-paired multimodal data for model training. Moreoever, the shared modality can be either sensor data or labels, as both can effectively describe the similarity of events. Instead of requiring synchronized multimodal data capturing the exact same event, we show that data of different modalities observing similar events, even captured at different times and locations, can be effectively used for multimodal training.
@@
Specifically, MMBind features a novel two-stage training strategy to bind distributed and heterogeneous IoT data for multimodal training, including (1) pairing incomplete data through shared modalities and (2) weighted contrastive learning with heterogeneous paired data. In the first stage, MMBind gathers all data samples containing the shared modality to pre-train a unimodal encoder. When the shared modality is the data label, pre-trained large language models (LLMs) can be leveraged instead. MMBind then utilizes the feature similarity of the shared modality to match the most similar samples among different sets of incomplete data, thereby constructing pseudo-paired multimodal data. MMBind can also be easily extended to integrate multiple incomplete datasets through successive data binding, even when utilizing different shared modalities. In the second stage, MMBind employs an adaptive multimodal learning architecture capable of training models on data with heterogeneous modality combinations. Such a design enables MMBind to leverage both pseudo-paired and original incomplete multimodal data for model training. Furthermore, MMBind performs weighted multimodal contrastive learning based on the similarity of data pairs, thereby accommodating different contributions from pseudo-paired samples to enhance the robustness of the learned representations. As a result, MMBind efficiently trains multimodal joint embeddings on large volumes of incomplete data from distributed nodes, requiring only limited or even no naturally paired multimodal data. Moreover, MMBind supports an efficient edge implementation framework. By aggregating incomplete data samples across edge devices to train the multimodal joint embeddings, individual edge devices can subsequently fine-tune the model for specific downstream tasks, resulting in low computational overhead.
@@
We extensively evaluate the performance of MMBind across ten different multimodal datasets, encompassing a total of nine distinct sensor modalities. The results indicate that MMBind achieves up to a 20% accuracy improvement over several state-of-the-art baselines in both cross-node and cross-dataset data binding, and reduces training overhead on edge devices compared to training models from scratch using complete multimodal data.
@@
  • Based on our key findings, we propose MMBind, a new data binding approach for multimodal learning with distributed and heterogeneous IoT data. MMBind leverages shared modalities to bind disparate and incomplete datasets, requiring limited or even no naturally paired data.
  • MMBind incorporates an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations, and a weighted contrastive learning approach to effectively handle domain shifts among disparate data sources.
@@ 2. Related Work
Distributed Sensing and Learning. Distributed sensor networks, consisting of spatially distributed sensing nodes, can vary from nodes in close proximity with overlapping fields of view (Samplawski et al., 2023; Großwindhager et al., 2019), to deployments in entirely different environments observing distinct scenes (Wu et al., 2022; Feng et al., 2021). Training models on such spatially distributed data often employs sensor fusion algorithms (Jeong et al., 2024; Chen et al., 2022; Li et al., 2021) that integrate information from multiple perspectives, or federated learning approaches (Ouyang et al., 2024, 2023) that aggregate models of different nodes. However, incorporating data from distributed nodes with heterogeneous sensor modalities for multimodal learning remains a significant challenge.
Multimodal Learning with Missing Modality. When performing multimodal learning in distributed multimodal sensing systems, one must contend with heterogeneous and incomplete data, where each sensor node captures a subset of the desired modalities. Several recent works have explored strategies for training with incomplete multimodal data. For example, (Liang et al., 2022; Huang et al., 2021) utilize unimodal encoders pre-trained on single-modal data to build a unified multimodal model, while (Wang et al., 2020; Cai et al., 2018) generate missing modality data from available modalities. However, these methods often struggle to effectively learn cross-modal information in IoT applications due to the scarcity of high-quality multimodal pairs. In contrast, MMBind leverages shared modalities across incomplete datasets to create paired data, allowing the model to learn cross-modal correlations more efficiently.
Model Binding Methods such as ImageBind. Recent works such as ImageBind (Girdhar et al., 2023) and LanguageBind (Zhu et al., 2023) align different modalities to the embedding space of one central modality through contrastive learning. (Dai et al., 2024b) expands upon this by sequentially training models of two modalities rather than relying on a central modality. All these approaches rely on model binding, where various modalities are indirectly aligned through transitivity, exposing the model only to (Central, X) modality pairs. In contrast, MMBind develops a data binding approach that explicitly synthesizes pairs of (Central, X, Y) data, enabling direct multimodal learning across multiple modalities. Additionally, MMBind utilizes the shared modality encoder only for data pairing, while ImageBind relies upon it to perform embedding space alignment across all modalities, resulting in greater susceptibility to insufficient shared modality data. Our evaluations in Section 6 further show the superiority of MMBind on IoT datasets with limited training data and significant domain shifts.
@@ 3. A Motivation Study
In this section, we evaluate multimodal learning performance with limited paired data and explore the potential of leveraging disparate and incomplete data using several real-world datasets. Here, “paired data” refers to full-modality data, with “pairing” describing the process of combining separate data samples into one with complete modalities. The key insights from these results motivate the design of MMBind.
@@
3.1. Impact of Limited Paired Data
Specifically, we evaluate using accelerometer, gyroscope and magnetic data from 15 subjects in the public RealWorld dataset (Sztyler and Stuckenschmidt, 2016) to classify five human activities (walking, sitting, standing, walking upstairs, and walking downstairs). To simulate varying amounts of paired multimodal data, we randomly divide the dataset into training sets containing 10%, 5%, and 1% of labeled data, and a testing set with 90%, while maintaining class balance. The deep learning model used consists of five CNN layers, two GRU layers, and one fully connected layer. Each experiment is repeated five times.
@@ 3.2. Potential of Binding Incomplete Data
Figure 3 shows that integrating pseudo-paired data from MotionSense and Shoaib enhances performance, especially when complete multimodal data is limited. Notably, pre-training the model with pseudo-paired samples significantly boosts the performance of fine-tuning with only 1% of training data (77.9%) from the RealWorld dataset, and even approaches the performance of 10% training data (79.1%).
@@ 4. System Overview
We now introduce MMBind, a new framework for multimodal learning with distributed and heterogeneous IoT data. We first introduce the applications and challenges, and then describe the problem formulation and system architecture.
@@
4.1. Applications and Challenges
MMBind is designed for a wide range of applications where heterogeneous sensors are deployed on distributed nodes to perform complex sensing tasks such as activity recognition (Ouyang et al., 2022), scene classification (Zeng et al., 2021) or vehicle recognition (Shi et al., 2022). In these applications, MMBind can harness heterogeneous and incomplete data collected by distributed nodes to effectively train a multimodal neural network.
We now describe two typical application scenarios. First, multimodal data across different nodes often exhibits significant heterogeneity, with missing modalities or labels. For example, sensors available on distributed nodes can change dynamically due to deployment constraints or runtime sensor failures. As illustrated in Figure 4, when the data is labeled, MMBind can pair samples from different nodes that describe the same or similar events. For instance, in activity recognition, one node may capture only IMU data while another records skeleton data. MMBind can pair IMU and skeleton data from the same or similar activity (e.g., walking and jogging) to train multimodal neural networks. When data labels are unavailable, MMBind can pair different incomplete datasets using shared sensor modalities. One node may collect IMU and audio data while another gathers IMU and skeleton data. MMBind can pair audio and skeleton data based on a computed similarity of their IMU data. Second, adapting multimodal neural networks to an expanded sensing suite is also challenging. For example, consider a substantial dataset of indoor scenes collected using devices with RGB cameras and depth sensors. If a new mmWave sensor is introduced, the old dataset will become incomplete in modality. In this scenario, MMBind can effectively make use of previously collected incomplete RGB and Depth data by pairing with the new mmWave data using shared modalities, reducing the amount of complete multimodal data to be recollected. In the following, we discuss these challenges in detail to motivate the design of MMBind.
Generating data pairs with heterogeneous data. The first challenge is how to generate effective full-modality data from different incomplete data samples. First, sensor data in IoT applications is usually highly heterogeneous, with varying dimensions and data patterns, making it nontrivial to quantify the similarity of the shared modality for data pairing. Second, when multiple overlapping modalities are present, selecting the optimal common modality becomes crucial for enhancing the effectiveness of data pairing.
@@ 4.3. System Architecture
The design of MMBind is motivated by the key insight from Section 3 that distributed data from different modalities, captured at various times and locations, can be leveraged to enhance multimodal training if they describe similar events. Therefore, our key idea is to bind incomplete multimodal data based on the shared modality, and create pseudo-paired multimodal data samples to learn multimodal joint embeddings. Figure 5 shows the overall system architecture of MMBind.
Specifically, MMBind features a new two-stage training strategy to bind distributed and heterogeneous IoT data for multimodal training, including pairing incomplete data with shared modalities and weighted contrastive learning with heterogeneous paired data. In the first stage, MMBind gathers all samples containing the shared modality and trains the corresponding unimodal encoder. This shared modality can be either sensor data or labels, as both can describe the similarity of events captured by distributed nodes. When the shared modality is a data label, pre-trained large language models (LLMs) can be leveraged to measure similarity of samples. Then, MMBind utilizes the feature similarity of the shared modality to construct pseudo-paired multimodal data samples by matching the most similar samples across different incomplete datasets. Finally, each pseudo-paired multimodal sample will be associated with a corresponding similarity measurement.
In the second stage, MMBind employs an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations, by representing missing modalities with dummy inputs. Therefore, it can integrate both pseudo-paired and original incomplete multimodal data into the training process. Furthermore, to account for imperfect pairing caused by domain shifts among disparate data sources, MMBind introduces weighted contrastive learning on the pseudo-paired multimodal data. MMBind assigns varying weights to samples in contrastive learning based on the similarity of data pairs, thereby accommodating different contributions from pseudo-paired samples. As a result, MMBind efficiently learns multimodal joint embeddings from distributed and heterogeneous IoT data, requiring only limited or even no naturally paired multimodal data.

### v2

@@ Abstract.
Multimodal sensing systems are increasingly prevalent in various real-world applications. Most existing multimodal learning approaches heavily rely on training with a large amount of synchronized, complete multimodal data. However, such a setting is impractical in real-world IoT sensing applications where data is typically collected by distributed nodes with heterogeneous data modalities, and is also rarely labeled. In this paper, we propose MMBind, a new data binding approach for multimodal learning on distributed and heterogeneous IoT data. The key idea of MMBind is to construct a pseudo-paired multimodal dataset for model training by binding data from disparate sources and incomplete modalities through a sufficiently descriptive shared modality. We also propose a weighted contrastive learning approach to handle domain shifts among disparate data, coupled with an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations. Evaluations on ten real-world multimodal datasets highlight that MMBind outperforms state-of-the-art baselines under varying degrees of data incompleteness and domain shift, and holds promise for advancing multimodal foundation model training in IoT applications [^1] The source code is available via https://github.com/nesl/multimodal-bind..
Multimodal Learning, Distributed IoT Data, Foundational Models
[…]
journalyear: 2025conference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 6–9, 2025; Irvine, CA, USAbooktitle: The 23rd ACM Conference on Embedded Networked Sensor Systems (SenSys ’25), May 6–9, 2025, Irvine, CA, USAdoi: 10.1145/3715014.3722053isbn: 979-8-4007-1479-5/25/05ccs: Human-centered computing Mobile computingccs: Computing methodologies Learning paradigms
@@ 1. Introduction
Multimodal sensing systems are increasingly deployed in real-world applications such as health monitoring (Ouyang et al., 2024; Zhang et al., 2020), Augmented Reality (AR) (Chen et al., 2018; Apicharttrisorn et al., 2022), localization (Jeong et al., 2024), and autonomous driving (Shi et al., 2022; He et al., 2023). Most existing multimodal learning approaches require extensive training with complete multimodal data, where each sample includes data from all sensor modalities collected simultaneously along with corresponding labels (Ouyang et al., 2022; Liu et al., 2024, 2021). However, this paradigm is impractical in real-world IoT sensing applications where data is typically collected by distributed nodes. First, such data is usually heterogeneous and incomplete—the available multimodal data on different nodes can vary significantly, often with missing modalities or labels. For example, sensor nodes deployed in homes may contain diverse sets of modalities due to deployment constraints or privacy concerns. Sensors on a specific node may also fail dynamically, resulting in changing sensor modalities at runtime. Labeling multimodal data in real-world settings is also challenging, as data from many sensors, such as IMU, are not intuitive for human annotation (Hu et al., 2020; Shuai et al., 2021). Second, sensor data collected by distributed nodes is typically gathered at different times and locations, resulting in disparate data that may describe similar events. For instance, in activity recognition, a wearable worn by an outdoor subject may only collect IMU data, while an indoor node may exclusively capture skeleton data of the same activity. Given the prevalence of disparate and heterogeneous data in IoT sensing, it is vital to effectively incorporate this data for multimodal learning.
@@
In this paper, we propose MMBind, a new framework for multimodal learning with distributed and heterogeneous IoT data. As shown in Figure 1, MMBind is the first data binding approach that leverages shared modalities across distributed nodes to creating pseudo-paired multimodal data for model training. Unlike existing model-binding approaches, which do not explicitly associate data modalities across different nodes, our data-binding approach demonstrates superior performance on IoT datasets with limited training data and significant domain shifts. Moreover, the shared modality can be either sensor data or labels, as both can effectively describe the similarity of events. Instead of requiring synchronized multimodal data capturing the exact same event, we show that data of different modalities observing similar events, even captured at different times and locations, can be effectively used for multimodal training. This enables MMBind to construct a multimodal foundational dataset from various disparate and incomplete small datasets collected by distributed IoT nodes, which is promising to advance multimodal foundation model training in IoT applications.
Figure 1. MMBind features a data binding approach that effectively integrates disparate data with heterogeneous modalities for multimodal learning, which outperforms the model binding approach, where various modalities are indirectly aligned through the encoder of a shared modality.
@@
Specifically, MMBind features a novel two-stage training strategy to bind distributed and heterogeneous IoT data for multimodal training, including (1) pairing incomplete data through shared modalities and (2) weighted contrastive learning with heterogeneous paired data. In the first stage, MMBind gathers all data samples containing the shared modality to pre-train a unimodal encoder. When the shared modality is the data label, pre-trained large language models (LLMs) can be leveraged instead. MMBind then utilizes the feature similarity of the shared modality to match the most similar samples from data with various domain shifts, thereby constructing pseudo-paired multimodal data that is associated with a corresponding similarity measurement. MMBind can also be easily extended to integrate multiple incomplete datasets through successive data binding, even when utilizing different shared modalities. In the second stage, MMBind employs weighted multimodal contrastive learning, assigning varying weights to samples during contrastive learning based on the similarity of data pairs. This design is unique to the data binding approach, as it accommodates different contributions from pseudo-paired samples. By considering the varying quality of pseudo-paired samples due to significant domain shifts, this method enhances the robustness of the learned representations. Furthermore, MMBind integrates an adaptive multimodal learning architecture capable of training models on data with heterogeneous modality combinations, leveraging both pseudo-paired and original incomplete multimodal data for model training. As a result, MMBind efficiently trains multimodal joint embeddings on large volumes of incomplete data from distributed nodes, requiring only limited or even no naturally paired multimodal data.
@@
We extensively evaluate the performance of MMBind across ten different multimodal datasets collected under uncontrolled and natural settings, encompassing a total of nine distinct sensor modalities and significant domain shifts across different subjects and environments. The results indicate that MMBind achieves up to a 20% accuracy improvement over several state-of-the-art baselines in both cross-node and cross-dataset data binding, and reduces training overhead on edge devices compared to training models from scratch using complete multimodal data.
@@
  • Based on our key findings, we propose MMBind, the first data binding approach for multimodal learning with distributed and heterogeneous IoT data. MMBind leverages shared modalities to bind disparate and incomplete datasets, requiring limited or even no naturally paired data.
  • MMBind incorporates a weighted contrastive learning approach based on the similarity of data pairs to effectively handle domain shifts among disparate data sources, and an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations.
@@ 2. Related Work
Distributed Sensing and Learning. Distributed sensor networks, consisting of spatially distributed sensing nodes, can vary from nodes in close proximity with overlapping fields of view (Samplawski et al., 2023; Großwindhager et al., 2019), to deployments in entirely different environments observing distinct scenes (Wu et al., 2022; Feng et al., 2021). Training models on such spatially distributed data often employs sensor fusion algorithms (Jeong et al., 2024; Chen et al., 2022; Li et al., 2021) that integrate information from multiple perspectives. Although previous multimodal federated learning studies (Ouyang et al., 2023; Wang et al., 2024) enable distributed model training over nodes with heterogeneous models, their upper bound is centralized unimodal learning, which does not fully align different modalities and has been shown to have inferior performance compared to MMBind in Section 6.
Multimodal Learning with Missing Modality. When performing multimodal learning in distributed sensing systems, one must contend with heterogeneous and incomplete data, where each sensor node captures a subset of the desired modalities. Several recent works have explored strategies for training with incomplete multimodal data. For example, (Liang et al., 2022; Huang et al., 2021) utilize unimodal encoders pre-trained on single-modal data to build a multimodal model, while (Wang et al., 2020; Cai et al., 2018) generate missing modality data from available modalities. However, these methods often struggle to effectively learn cross-modal information in IoT applications due to the scarcity of high-quality multimodal pairs. In contrast, MMBind leverages shared modalities across incomplete datasets to create paired data, allowing the model to learn cross-modal correlations more efficiently.
Model Binding Methods such as ImageBind. Recent works such as ImageBind (Girdhar et al., 2023) and LanguageBind (Zhu et al., 2023) align different modalities to the embedding space of one central modality through contrastive learning. (Dai et al., 2024b) expands upon this by sequentially training models of two modalities rather than relying on a central modality. All these approaches rely on model binding, where various modalities are indirectly aligned through transitivity, exposing the model only to (Central, X) modality pairs. In contrast, MMBind develops the first data binding approach that explicitly synthesizes pairs of (Central, X, Y) data based on the shared modalities, enabling direct multimodal learning across multiple modalities. Additionally, MMBind utilizes the shared modality encoder only for data pairing, while ImageBind relies upon it to perform embedding space alignment across all modalities, resulting in greater susceptibility to insufficient shared modality data. Our evaluations in Section 6 further show the superiority of MMBind on IoT datasets with limited training data and significant domain shifts.
@@ 3. A Motivation Study
In this section, we evaluate multimodal learning performance with limited paired data and explore the potential of leveraging disparate and incomplete data. Here, “paired data” refers to full-modality data, with “pairing” describing the process of combining separate data samples into one with complete modalities.
3.1. Impact of Limited Paired Data
Specifically, we evaluate using data from 15 subjects in the public RealWorld dataset (Sztyler and Stuckenschmidt, 2016) to classify five human activities (walking, sitting, standing, walking upstairs, and walking downstairs). To simulate varying amounts of paired multimodal data, we divide the dataset into training sets containing 10%, 5%, and 1% of labeled data, and a testing set with 90%, while maintaining class balance. The deep learning model used consists of five CNN layers, two GRU layers, and one fully connected layer. We repeat each experiment on five different seeds.
@@ 3.2. Potential of Binding Incomplete Data
Figure 3 shows that integrating pseudo-paired data from MotionSense and Shoaib enhances performance, especially when complete multimodal data is limited. Notably, pre-training the model with pseudo-paired samples significantly boosts the performance of fine-tuning with only 1% of training data (77.9%), and even approaches the performance of 10% training data (79.1%).
@@ 4. System Overview
We now introduce MMBind, a new framework for multimodal learning with distributed and heterogeneous IoT data. The key idea is to bind disparate data to create pseudo-paired multimodal data for model training. We first introduce applications and challenges, and then describe the problem formulation and system architecture.
4.1. Applications and Challenges
MMBind is designed for a wide range of applications where heterogeneous sensors are deployed on distributed nodes to perform complex sensing tasks such as activity recognition (Ouyang et al., 2022), scene classification (Zeng et al., 2021) or vehicle recognition (Shi et al., 2022). In these applications, MMBind can harness heterogeneous and incomplete data collected by distributed nodes to effectively train a multimodal neural network. MMBind can also construct a multimodal foundational dataset from various incomplete small datasets, which is promising to advance multimodal foundation model training in IoT applications.
We now describe two typical application scenarios. First, multimodal data across different nodes often exhibits significant heterogeneity, with missing modalities or labels. As illustrated in Figure 4, when the data is labeled, MMBind can pair samples from different nodes that describe the same or similar events. For instance, in activity recognition, one node may capture only IMU data while another records skeleton data. MMBind can pair IMU and skeleton data from the same or similar activity (e.g., walking and jogging) to train multimodal neural networks. When data labels are unavailable, MMBind can pair different incomplete datasets using shared sensor modalities. One node may collect IMU and audio data while another gathers IMU and skeleton data. MMBind can pair audio and skeleton data based on a computed similarity of their IMU data. Second, adapting multimodal neural networks to an expanded sensing suite is also challenging. For example, consider a substantial dataset of indoor scenes collected using devices with RGB cameras and depth sensors. If a new mmWave sensor is introduced, the old dataset will become incomplete in modality. In this scenario, MMBind can effectively make use of previously collected incomplete RGB and Depth data by pairing with the new mmWave data using shared modalities, reducing the amount of complete multimodal data to be recollected. In the following, we discuss these challenges in detail to motivate the design of MMBind.
Generating data pairs with heterogeneous data. The first challenge is how to generate effective full-modality data from incomplete data samples. First, sensor data in IoT applications is usually highly heterogeneous, with varying dimensions and data patterns, making it nontrivial to quantify the similarity of the shared modality for data pairing. Second, when multiple overlapping modalities are present, selecting the optimal common modality becomes crucial for enhancing the effectiveness of data pairing.
@@ 4.3. System Architecture
The design of MMBind is motivated by the key insight from Section 3 that distributed data from different modalities, captured at various times and locations, can be leveraged to enhance multimodal training if they describe similar events. Therefore, our key idea is to bind incomplete multimodal data based on the shared modality, and create pseudo-paired multimodal data samples to learn multimodal joint embeddings. Figure 5 shows the overall system architecture.
Specifically, MMBind features a new two-stage training strategy to bind distributed and heterogeneous IoT data for multimodal training, including pairing incomplete data with shared modalities and weighted contrastive learning with heterogeneous paired data. In the first stage, MMBind gathers all samples containing the shared modality and trains the corresponding unimodal encoder. This shared modality can be either sensor data or labels, as both can describe the similarity of events captured by distributed nodes. When the shared modality is a data label, pre-trained large language models (LLMs) can be leveraged to measure similarity of samples. Then, MMBind utilizes the feature similarity of the shared modality to construct pseudo-paired multimodal data samples by matching the most similar samples across different incomplete datasets. Finally, each pseudo-paired multimodal sample will be associated with a corresponding similarity measurement. In the second stage, MMBind employs an adaptive multimodal learning architecture capable of training models with heterogeneous modality combinations, by representing missing modalities with dummy inputs. Therefore, it can integrate both pseudo-paired and original incomplete multimodal data into the training process. Furthermore, to account for imperfect pairing caused by domain shifts among disparate data sources, MMBind introduces weighted contrastive learning on the pseudo-paired multimodal data. MMBind assigns varying weights to samples in contrastive learning based on the similarity of data pairs, thereby accommodating different contributions from pseudo-paired samples. As a result, MMBind efficiently learns multimodal joint embeddings from distributed and heterogeneous IoT data, requiring only limited or even no naturally paired multimodal data.